# Enhancing Wide-Angle Image Using Narrow-Angle View of the Same Scene

### v1

@@ Abstract
A common dilemma while photographing a scene is whether to capture it in wider angle, allowing more of the scene to be covered but in lesser details or to click in narrow angle that captures better details but leaves out portions of the scene. We propose a novel method in this paper that infuses wider shots with finer quality details that is usually associated with an image captured by the primary lens by capturing the same scene using both narrow and wide field of view (FoV) lenses. We do so by training a GAN-based model to learn to extract the visual quality parameters from a narrow angle shot and to transfer these to the corresponding wide-angle image of the scene. We have mentioned in details the proposed technique to isolate the visual essence of an image and to transfer it into another image. We have also elaborately discussed our implementation details and have presented the results of evaluation over several benchmark datasets and comparisons with contemporary advancements in the field.
@@
Ultra-wide Image , Super Resolution , Generative Adversarial Network
[…]
[auth1]organization=Department of Computer Science and Engineering, University of Dhaka, city=Dhaka
[…]
[auth2]organization=Department of Computer Science and Engineering, Jahangirnagar University, city=Dhaka
[…]
[auth3]organization=Department of Computer Science and Engineering, University of Dhaka, city=Dhaka
@@ 1 Introduction
Ultra-wide photography is a popular photographic trend in today’s world. This form of image-taking, though accommodating a greater field of view and thus offering a unique aesthetic, degrades the finer details better retained by the primary lens. This trade-off between the field of view covered and the quality of the finer details retained by the captured image places the users in an unwanted dilemma where one has to compromise the field of view for the details or vice-versa. Our primary goal in this paper is to enable the camera users to capture wide-angle images with quality that is identical to the narrow-angle shots while covering the original wide field of view.
The fundamental idea of capturing wide-angle images that retain the finer details associated with narrow-angle shots is to capture the scene of interest with two different lenses, both narrow and wide. The expectation here is that the narrow-angle shot provides the expected resolution while the wide-angle one carries the contents of the scene not covered in the narrow-angle counterpart. The idea would then be to render the contents captured by the wide-angle lens in the resolution that is represented by the narrow-angle counterpart. As it has been shown by the paper on Neural Style Transfer [6], the style and content of an image are separable and the stylistic representation of an image can be transferred into another image. We have leveraged this finding to transfer the resolution qualities of the narrow-angle shot into the wide-angle counterpart. Furthermore, we have implemented a pipeline that incrementally transfers the qualitative representations from the narrow-angle shot to the immediate wide-angle shot and repeats the process until the widest available image (often 0.5x or ultra-wide image) is reached. This involves capturing the scene of interest using all available lenses and incrementally applying the finest available quality (from the image with the narrowest field of view) through the stack of available images up to the image with the widest field of view.
In this paper, we propose a GAN-based architecture where the generator network produces realistic-looking enhanced versions of the wide FoV image while the discriminator component attempts to identify generated high-resolution results from the ground truth. In the first step, the spatial resolution of the original wide FoV image is upscaled using bicubic interpolation so that more details can be physically accommodated. Once upscaled, both the enhanced wide (from now onwards to be referred to as the wide FoV) and the narrow FoV images are partitioned into smaller patches so that visually similar patches may be utilized to assist the process of resolution transfer, inspired by the paper on Vision Transformer [11]. For the purpose of resolution transfer, we propose a modified version of the Gram matrix representation of an image, as used in the paper on Neural Style Transfer. Upon transferring the resolution from the narrow FoV patches into that from the wide FoV image, the synthetic image is passed through the discriminator setup that refines the seams between the tiles and the overall perceptual fidelity of the image, leveraging the task-specific loss functions we have designed. Finally, we perform this routine of resolution transfer throughout the stack of all available lenses to attain the highest possible resolution for the widest available field of view. Our contributions can thus be summarized into the following points,
  • 1.
    An end-to-end pipeline capable of transferring finer quality details to the wide-angle image from the corresponding narrow-angle image
  • 2.
    The novel concept of transferring visual quality across the stack of all available lenses
  • 3.
    An innovative modification of the Gram matrix that is able to capture the visual essence of an image
@@ 2 Related Works
Enhancing image quality has remained a central pursuit in computer vision. Earlier methods were primarily centered around super-resolution or upscaling techniques [1], which aimed to reconstruct high-resolution images from their low-resolution counterparts. While effective in producing sharper visuals, these models often introduced artifacts like oversmoothing, leading to a loss in texture and fine details. Later, generative adversarial models such as SRGAN [3], ESRGAN [2], and Real-ESRGAN [4] improved perceptual realism by emphasizing texture fidelity and producing visually convincing outputs. However, these models still struggle with preserving structural consistency, especially in scenes where the narrow and wide FOV images exhibit differences in perspective or content.
@@
Recent advancements have introduced lightweight and efficient architectures tailored for real-time applications. LatticeNet [24] employs a lattice block structure inspired by filter banks, enabling adaptive feature fusion while maintaining low computational complexity. This design facilitates the reconstruction of high-quality images without the overhead associated with deeper networks.
Building upon the success of attention mechanisms, the Efficient Long-Range Attention Network (ELAN) [25] integrates shift convolutions with group-wise multi-scale self-attention modules. This combination captures both local and long-range dependencies, enhancing the model’s ability to reconstruct fine details across varying scales. ELAN’s architecture achieves a balance between performance and efficiency, making it suitable for applications requiring real-time processing. Further exploring attention-based models, DLGSANet [23] introduces a hybrid approach combining dynamic local and sparse global self-attention mechanisms. The Multi-Head Dynamic Local Self-Attention (MHDLSA) module effectively captures local features, while the Sparse Global Self-Attention (SparseGSA) module focuses on the most relevant global information. This dual attention strategy enhances the model’s capacity to reconstruct high-resolution images with reduced computational demands. In parallel, the Deep Multi-level Feature Fusion Network (DMFFN) [26] addresses common challenges in super-resolution tasks, such as blurred edges and redundant structures. By employing two-level nested residual blocks with asymmetric structures, DMFFN efficiently fuses multi-level features, leading to improved edge preservation and faster convergence during training.
These contemporary models demonstrate significant progress in balancing performance and efficiency, particularly in scenarios requiring the enhancement of wide FOV images by leveraging information from narrow FOV counterparts. Their architectures are designed to maintain structural consistency and fine details, addressing the limitations observed in earlier methods.
Moreover, the integration of perceptual metrics such as LPIPS [16] has become increasingly prevalent, offering a more nuanced assessment of visual similarity that aligns closely with human perception. Incorporating human-in-the-loop evaluations further ensures that the enhanced images meet real-world expectations for quality and accuracy.
In this work, building upon these advancements, we aim to improve wide FOV image quality by transferring visual fidelity from narrow FOV versions of the same scene. By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, we target high-quality, structurally accurate image synthesis suited for immersive visual applications.
@@ 3 Method
In this section, we have described in detail the adopted methodologies and architecture of the model used to transfer high-quality details from the narrow-angle image into the wide-angle counterpart.
Figure 1: Model Architecture
In the case of training the model, we consider the original image as the wide FoV ground truth and extract the central portion to serve as the high-definition narrow FoV. The original wide FoV is then downsampled to simulate the captured wide FoV image.
3.1 Content Reconstruction
Our primary goal is to reconstruct the contents of the wide-angle image applying the visual qualities of the narrow-angle counterpart. By visual qualities we indicate the parameters that impact the perceptual cues which are color, brightness, contrast, and sharpness. For reconstructing the contents of the wide-angle image, we utilize the pixel-wise content loss. We predict the pixel entries for a white-noise image such that the following content loss is minimized.
$L^{(l)}_{content}(p,q)=\frac{1}{2}\sum_{i,j}(Q^{l}_{ij}-P^{l}_{ij})^{2}$
Here $p$ and $q$ are the original and white-noise images respectively and $P^{l}_{ij}$ and $Q^{l}_{ij}$ are the feature representations in the $l^{th}$ layer for $i^{th}$ filter at the position $j$.
3.2 Visual Quality Reconstruction
As mentioned earlier, the visual qualities are obtained from the narrow-angle image. The qualities are extracted from the image using the gram matrix approach. We have divided the image into a fixed number of patches and computed Gram matrices for each patch. The motivation to adopt this approach comes from the intuition that visually similar patches may exist between the narrow and the wide-angle images. Obtaining visual qualities from the patch of narrow-angle image that visually aligns with a patch of the wide-angle image would produce better results. Hence, we divide the entire narrow and wide-angle images into a predefined number of patches. For each patch of the wide-angle image, the patch from the narrow-angle image that is visually most similar is selected. The visual similarity score between patches is determined using a perceptual loss function described in the following section. Patches are taken to be visually similar if the similarity score exceeds a threshold. In case no patch from the narrow-angle image is found to exceed the threshold, a gram matrix representation of the entire narrow-angle image is used as a backup.
Figure 2 illustrates partitioning the image into fixed-sized patches. Images partitioned into fixed-sized patches just for illustrative purposes do not indicate the actual number of patches or their proportions used in the implementation. The patches are bound by black lines. Arrows and bounding boxes of different colors denote patches in the narrow FoV image that are visually similar to certain patches from the wide FoV image.
@@ 3.2 Visual Quality Reconstruction
3.2.1 Dividing into Patches
The narrow and wide-angle images are first divided into patches of dimension $128\times 128$. This particular size for the patches has been chosen to create a balance between structural information contained within an image and the computational cost incurred to process the patches. The idea of dividing the images into patches also takes inspiration from the paper on Vision Transformer [11].
3.2.2 Patch-wise Similarity
We use perceptual loss [17] to quantify the visual similarity between patches from narrow and wide-angle images. We use the mid layers of a VGG19 network to numerically determine the perceptual loss between the pair of patches. Following function is used to determine the loss,
$L_{feature}(p_{1},p_{2})=\frac{1}{C_{j}*H_{j}*W_{j}}||\phi_{j}(p_{2})-\phi_{j}(p_{1})||_{2}^{2}$
Here $p_{1}$ and $p_{2}$ are the patches from wide and narrow-angle images, $\phi_{j}(p_{1})$ and $\phi_{j}(p_{2})$ denote the feature map obtained from the $j^{th}$ layer, and $C_{j},H_{j}$ and $W_{j}$ are the number of channels, height, and width of the patches.
3.3 Generator Architecture
Figure 3: Generator Architecture
We utilize a ResNet [27] based generator with residual blocks. The input layer is followed by a collection of residual blocks to facilitate earlier layers in learning complex functions. The residual block relays the data into the upsampling block that generates the enhanced patch and passes through a final convolution layer to adjust the channel dimension.
3.3.1 Feature Map Construction
We consider the following parameters to effectively represent the visual qualities of an image (patch of an image in our case); $color,contrast,brightness,sharpness$. Considering each of these parameters as distinct features of a patch, an enhanced feature map representation of the patch is computed. The enhanced feature map is an augmentation of the individual features extracted from the patch and is formulated as follows,
$\hat{F_{l}}=[F_{l},F_{l_{color}},F_{l_{brightness}},F_{l_{contrast}},F_{l_{sharpness}}]$
Where $F_{l}\epsilon\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the feature map extracted using a VGG backbone network. Rest of the features are computed as follows,
@@ 3.3.1 Feature Map Construction
    Color: An average of the RGB channels forms an embedding that represents the overall intensities. Computed as,
    $F_{l_{color}}=\frac{1}{H_{l}W_{l}}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{j=1}F_{l}$
@@
    Brightness: Mean pixel intensities over all pixels are taken to indicate the perceived luminance of the patch. Computed as,
    $F_{l_{brightness}}=\frac{1}{C_{l}H_{l}W_{l}}\sum^{C_{l}}_{i=1}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{k=1}F_{l}$
@@
    Contrast: Variance of the pixel intensities is computed to represent the variation of intensities, that is the contrast, within the patch. Calculated as,
    $F_{l_{contrast}}=\sqrt{\frac{1}{C_{l}H_{l}W_{l}}\sum^{C_{l}}_{i=1}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{k=1}(F_{l}-\mu_{l})^{2}}$
    here $\mu_{l}$ is the average intensity.
  4. 4.
    Sharpness: Sharpness is captured using the Laplacian filter. The Laplacian filter is multiplied to the feature map as follows,
    $F_{l_{sharpness}}=F_{l}*\bigtriangledown^{2}$
3.3.2 Gram Matrix Computation
Once the feature map is constructed, we can use this representation to find the gram matrix representation of the patch of the image for the current layer. The computation can be performed as follows,
$G_{l}=\hat{F}_{l}\hat{F}_{l}^{\top}$
The gram matrix is the product of the enhanced feature map to its transpose, essentially capturing the pair-wise correlation between features.
3.3.3 Loss Calculation
Visual quality loss is defined as the squared distance between the gram matrix representation of the source and target images. In our case, the source is the patch from the narrow FoV image while the target is a white-noise image which is eventually going to ”be” the resultant image by minimizing the loss defined above. The numerical representation of the loss is,
$L^{(l)}_{visual}=\sum_{l}w_{l}||G_{l_{q}}-G_{l_{p}}||^{2}$
Here $G_{l_{q}}$ and $G_{l_{p}}$ are the gram matrix representations of the patches from the white-noise and the narrow FoV images respectively. Mentionable that in case no particular match is found for a certain wide FoV patch in the narrow FoV, $G_{l_{p}}$ would denote the gram matrix representation of the entire narrow FoV image.
Hence the total generator loss is quantified as,

### v2

@@ Abstract
A common dilemma while photographing a scene is whether to capture it at a wider angle, allowing more of the scene to be covered but in less detail or to click in a narrow angle that captures better details but leaves out portions of the scene. We propose a novel method in this paper that infuses wider shots with finer quality details that is usually associated with an image captured by the primary lens by capturing the same scene using both narrow and wide field of view (FoV) lenses. We do so by training a Generative Adversarial Network (GAN)-based model to learn to extract the visual quality parameters from a narrow-angle shot and to transfer these to the corresponding wide-angle image of the scene using residual connections and an attention-based fusion module. We have mentioned in details the proposed technique to isolate the visual essence of an image and to transfer it into another image. We have also elaborately discussed our implementation details and have presented the results of evaluation over several benchmark datasets and comparisons with contemporary advancements in the field.
@@
Ultra-wide Image, Super Resolution, Generative Adversarial Network, Attention-based Fusion
@@ 1 Introduction
Modern smartphones and mirrorless cameras increasingly incorporate ultra-wide lenses [1] to capture expansive scenes in a single frame. While these lenses deliver a striking field of view, they often struggle to preserve the fine-grained detail achieved by conventional “primary” lenses, as discussed in the paper [2]. Users face a perennial trade-off: opting for a wider perspective at the expense of sharpness or retaining detail by narrowing the view. In this work, we seek to resolve this dilemma by synthesizing wide-angle images whose perceptual quality matches that of narrow-angle shots, without sacrificing the original field of view.
Our approach hinges on jointly leveraging both narrow- and wide-angle captures of the same scene. The narrow-angle image supplies high-resolution texture and detail, while the wide-angle image carries context beyond the primary lens’s range. We introduce an end-to-end pipeline that incrementally transfers visual fidelity from the finest (narrowest) view through to the widest view in a multi-lens stack. First, we partition both the wide image and the corresponding narrow-angle image into spatially aligned patches inspired by the paper on Vision Transformer [3]. Each patch pair is fed into a generator, which fuses content and detail via cross-view attention and further super-resolves the wide patch using pixel-shuffle layers—progressively enhancing resolution and texture across the full frame.
Central to our method is a novel GAN [4] architecture: a generator tailored to fuse cross-view content and detail features through attention mechanisms and a discriminator designed to enforce both global realism and seamless patch blending. We further extend the classic Gram matrix representation to encode not only texture but also color, brightness, contrast, and sharpness metrics, ensuring rich visual fidelity. Applying this procedure iteratively across all available zoom levels—from telephoto to ultra-wide—we end up with a single image that combines the sweeping perspective of the widest lens with the crisp detail of the primary lens. Our key contributions are
  • An end-to-end pipeline that transfers high-resolution detail from narrow to wide-angle images.
  • A GAN framework with cross-view attention and an enhanced Gram matrix formulation for robust detail propagation.
  • A cascading lens-stack strategy that scales seamlessly from telephoto to ultra-wide captures.
@@ 2 Related Works
Enhancing image quality has remained a central pursuit in computer vision. Earlier methods were primarily centered around super-resolution or upscaling techniques [5], which aimed to reconstruct high-resolution images from their low-resolution counterparts. While effective in producing sharper visuals, these models often introduced artifacts like oversmoothing, leading to a loss in texture and fine details. Later, generative adversarial models such as SRGAN [6], ESRGAN [7], and Real-ESRGAN [8] improved perceptual realism by emphasizing texture fidelity and producing visually convincing outputs. However, these models still struggle with preserving structural consistency.
@@
Figure 1: Overall model architecture
Recent advancements have introduced lightweight and efficient architectures tailored for real-time applications. LatticeNet [12] employs a lattice block structure inspired by filter banks, enabling adaptive feature fusion while maintaining low computational complexity. This design facilitates the reconstruction of high-quality images without the overhead associated with deeper networks. Building upon the success of attention mechanisms, the Efficient Long-Range Attention Network (ELAN) [13] integrates shift convolutions with group-wise multiscale self-attention modules. This combination captures both local and long-range dependencies, enhancing the model’s ability to reconstruct fine details across varying scales. The ELAN architecture achieves a balance between performance and efficiency, making it suitable for applications that require real-time processing.
Further exploring attention-based models, DLGSANet [14] introduces a hybrid approach combining dynamic local and sparse global self-attention mechanisms. The Multi-Head Dynamic Local Self-Attention (MHDLSA) module effectively captures local features, while the Sparse Global Self-Attention (SparseGSA) module focuses on the most relevant global information. This dual attention strategy enhances the model’s capacity to reconstruct high-resolution images with reduced computational demands. In parallel, the Deep Multilevel Feature Fusion Network (DMFFN) [15] addresses common challenges in super-resolution tasks, such as blurred edges and redundant structures. Using nested two-level residual blocks with asymmetric structures, DMFFN efficiently fuses multilevel features, leading to improved edge preservation and faster convergence during training. These contemporary models demonstrate significant progress in balancing performance and efficiency. Their architectures are designed to maintain structural consistency and fine details, addressing the limitations observed in earlier methods.
Perceptual metrics such as LPIPS [16] have become widely adopted for evaluating visual similarity in a way that aligns closely with human perception. Beyond perceptual loss, incorporating visual cues that guide perception—namely color, brightness, contrast, and sharpness—is equally important. We select these attributes based on extensive literature: [17] identifies them as key appearance properties in image quality assessment; [18] highlights the role of color in quality estimation; [19] emphasizes the critical balance of brightness and contrast in enhancement; and [20] combines contrast and sharpness to assess enhancement reliability.
By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, we target high-quality, structurally accurate image synthesis suited for immersive visual applications.
@@ 3 Method
The core philosophy of our solution is to use a cross-view attention mechanism to fuse content and visual quality features from a given pair of wide and narrow-angle images. The images are initially partitioned into fixed-size patches. We pick a pair of wide and narrow patches from the pool and pass them across the network architecture, as described below.
3.1 Image Partitioning
Both narrow and wide FoV images are partitioned into $128\times 128$-sized patches to ensure portions of images sharing visual similarity are used to exchange content and visual quality data. Once partitioned, patches from wide and narrow FoV images that share visual properties are mapped using perceptual loss [21] and stored in the pool, ready to be pushed into the generator network. A cosine similarity on the feature maps generated by the mid layers of a VGG-19 [22] backbone is used to determine the perceptual distance between patches.
@@ 3.1 Image Partitioning
Figure 3: Detailed architecture for the Generator component
3.2 Network Architecture
The architecture of the generator network is built upon a stack of ResNet-like [23] residual blocks that take as inputs a pair of wide and narrow FoV patches from the pool. The patches are encoded using separate encoder components to capture structural (content) and visual features from wide and narrow FoV patches respectively. Each block then employs a cross-view attention mechanism to fuse the content and visual properties from wide and narrow feature maps respectively. The pipeline consists of the following major components,
3.2.1 Structural Feature Encoder
Wide FoV patches are encoded with traditional CNN-based encoders to capture the content details. Encoded wide patch represented as $E_{w}$
3.2.2 Visual Feature Encoder
These encoders are specialized in extracting the visual parameters from a patch and are used to encode the narrow FoV patches. We use the Gram matrix approach to represent the visual features of a patch and include the following additional parameters in the Gram matrix representation: color, brightness, contrast, and sharpness. The feature map is constructed as follows,
  • The original feature map $F_{l}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is generated using a VGG-19 network.
  • Color information for each channel over spatial dimensions, also referred to as global average pooling [24], is computed as:
    $[F_{l_{\text{color}}}]_{c}$ $=\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w),$
    for $\quad\text{for }c=1,\dots,C_{l}$
  • Global mean pixel intensity across all channels and spatial dimensions indicates the perceived luminance (brightness) of the patch. Computed as:
    $F_{l_{\text{brightness}}}=\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w)$
  • We represent the overall contrast of the patch using the standard deviation:
    $F_{l_{\text{contrast}}}=\sqrt{\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}\left(F_{l}(c,h,w)-\mu_{l}\right)^{2}}$
    where $\mu_{l}=F_{l_{\text{brightness}}}$ is the global mean defined above.
  • Sharpness is accounted for using the Laplacian operated, as mentioned in this paper [25]:
    $F_{l_{\text{sharpness}}}(c,h,w)=(F_{l}(c,\cdot,\cdot)*\nabla^{2})(h,w)$
    where * denotes the convolution and $\nabla^{2}$ is the Laplacian kernel.
  • Finally, we augment the original feature map obtained from VGG-19 as follows:
    $\hat{F}_{l}=[F_{l},F_{l_{\text{color}}},F_{l_{\text{brightness}}},F_{l_{\text{contrast}}},F_{l_{\text{sharpness}}}]$
    where scalar or vector features are broadcast to match the spatial dimensions of $F_{l}$.
Once the enhanced feature map has been computed, we perform a dot product between the enhanced feature map and its transpose to attain the Gram matrix representation $E_{n}$, carrying the visual essence of the narrow FoV patch.
$E_{n}=\hat{F}_{l}\hat{F}_{l}^{\top}$
3.2.3 Cross-Attention Residual Blocks
Once the patches are encoded into the required latent space, we employ an attention-based fusion mechanism that fetches the most relevant visual features from the narrow patch and fuses those with the content features generated from the wide patch.
  • Encoded wide FoV patch $E_{w}$, representing the content data, acts as the query matrix
  • Encoded narrow FoV patch $E_{n}$, carrying the visual properties, provides both the key and value matrices
  • The attention scores, which gather semantically aligned, superior quality visual features present in the encoded narrow FoV patch, are computed using the scaled dot product equation introduced in the attention paper [26],
    $Attention(E_{n},E_{w})=softmax(\frac{E_{w}E_{n}^{\top}}{\sqrt{d_{k}}})E_{n}$
  • The gathered attention scores are then incorporated into the wide FoV feature map using a residual addition, producing the feature map representation of the enhanced wide FoV patch, $E$
    $E=E_{w}+Attention(E_{n},E_{w})$
3.2.4 Upsampling Layers
The output of the final residual block is passed through a pixel shuffle [27] layer with a scale factor equal to 2 to super-resolve the patch, denoted by $E^{*}$.
3.3 Loss Calculation
Generator loss is calculated patch-wise, i.e., each of the generated patches is compared against both the original wide and narrow FoV patches to quantify the content and visual deviations from the ground truth. Then the patches are assembled using seamless cloning, to eliminate visible seams during blending, and the reconstructed image is passed through the discriminator which computes a seam consistency loss over the entire image. The losses are described as follows,
3.3.1 Generator Loss
@@
    Content Loss: We find the deviations of the contents of the generated patch and ground truth with the following pixel-wise difference,
    $L^{(l)}_{content}(E^{o}_{w},E^{g}_{w})=\frac{1}{n}\sum_{i}(E^{o}_{w}-E^{g}_{w})^{2}$
    Here $E^{o}_{w}$ and $E^{g}_{w}$ are the latent space representations of the original and generated patches.
@@
    Visual Quality Loss: This loss is calculated as the difference of the Gram matrix representation of the generated patch and the narrow FoV patch using the following equation,
    $L^{(l)}_{visual}=\sum_{l}w_{l}||E^{o}_{n}-E^{g}_{n}||^{2}$
    Here $E^{g}_{n}$ and $E^{o}_{n}$ are the gram matrix representations of the patches from the generated and the original narrow FoV patches, respectively.
@@
    Total generator loss can be written as,